INAF queue observing, 2015 Jan 10-11

Observer: E. Sani, S. Bisogni
Telescope Operator: G. Bechetti
Support Astronomer: Michelle Edwards
Instrument: MODS, LUCI

Summary

The dome stays closed due to humidity, the temperature reaches the dew point

Data Summary

Project_RUN Objects Filter ExpTime Notes
MODS calibration
LUCI calibration

INAF queue observing, 2015 Jan 11-12

Observer: E. Sani, S. Bisogni
Telescope Operator: G. Bechetti
Support Astronomer: Michelle Edwards
Instrument: MODS, LUCI

Summary

At the beginning of the night the dome stays closed due to humidity, the temperature reaches the dew point. Humidity 100%, dewpoint -2.4, external temperature -2.6. No wind.

Data Summary

Project_RUN Objects Filter ExpTime Notes
LUCI Calibrations

INAF queue observing, 2015 Jan 12-13

Observer: E. Sani, S. Bisogni
Telescope Operator: G. Bechetti
Support Astronomer: Dave Thompson (on call)
Instrument: LUCI

Summary

At the beginning of the night the dome stays closed due to humidity.

Data Summary

Project_RUN Objects Filter ExpTime Notes
LUCI Calibrations

INAF queue observing, 2015 Jan 13-14

Observer: E. Sani, S. Bisogni
Telescope Operator: D. Gonzales Huerta
Support Astronomer: Dave Thompson (on call)
Instrument: ---

Summary

At the beginning of the night the dome stays closed due to humidity and strong wind.

Data Summary

Project_RUN Objects Filter ExpTime Notes

INAF queue observing, 2015 Jan 14-15

Observer: E. Sani, S. Bisogni
Telescope Operator: D. Gonzales Huerta
Support Astronomer: Michelle Edwards (on call)
Instrument: LUCI-MODS-LBC

Summary

During the twilight the dome is closed due to high humidity (>95%) and temperature clore to the dew point (-8.1C). Thin cirrus are passing, clouds on the NE horizon.
Once the dome is opened we have thin cirrus scattered over the sky.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Antoniucci hip27881 (vytau standard) grat 200H+K filters HK - zJ 4s  
MIgnoli Feige34 red grat only filter s_SDSS 6min  
MIgnoli MOS22 red grat only filter s_SDSS 1.5 h  
MIgnoli LBC_red_MOS22 filter z_SDSS    

INAF queue observing, 2015 Jan 15-16

Observer: E. Sani, S. Bisogni
Telescope Operator: D. Gonzales Huerta
Support Astronomer: Michelle Edwards (on call) and Barry Rothberg (at the telescope)
Instrument: MODS - LBCred

Summary

During the twilight the sky is mostly clear, but we have stron wind with maximum gusts at 24.5 m/s. We can't run skyflats with LBCred.
At the dome opening we have cirrus scattered over the south horizon, maximum wind gusts 17.5 m/s and low humidity (16%).
Extremely variable seeing during the night ranging from 0.7 to 1.8.

Data Summary

Project_RUN Objects Filter ExpTime Notes
SaraccoB RDCS red gray only filter r_SDSS 2h variable seeing
MIgnoli MOS22 red grat only filter r_SDSS 2 h pseudomonucular mode. variable seeing
MIgnoli LBC_red_MOS22 filter z_SDSS 30m  
MIgnoli LBC_red_MOS22 filter i_SDSS 30m  
MIgnoli LBC_red_MOS22 filter r_SDSS 30m  
MIgnoli LBC_red_MOS22 filter Y-FAN 12m  
MIgnoli LBC_red SA104 standard filter r_, i_, z_SDSS, Y-FAN   airmass 1.15

-- %USERSIG{RobertaCarini - 2015-02-09}%

Comments

 
Topic revision: r5 - 08 Jan 2016, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback